OpenStudy (anonymous):

write an algebraic expprestion that represents the perimeter of the rectangle below the bottom is 2x+1 the side is 5-3x.

OpenStudy (anonymous):

Perimeter is 2*side + 2*bottom \[2(2x+1)+2(5-3x)=4x+2+10-6x=12-2x\]

OpenStudy (anonymous):

what is the answer?

OpenStudy (anonymous):

The answer is at the end after the equals sign.... \[12-2x\]